OpenStudy (anonymous):

(16x3 + 2x2) - 4x3 =

OpenStudy (anonymous):

12x^3+2x^2=0 x^2(12x+2)=0 x=0,0,-1/6

OpenStudy (anonymous):

you know the difference?

OpenStudy (anonymous):

What difference?

OpenStudy (lgbasallote):

why two 0? lol

OpenStudy (anonymous):

Find the difference. (16x3 + 2x2) - 4x3 = A. 14x3 B. 12x3 - 2x2 C. 12x3 + 2x2 D. 16x3 - 2x2

OpenStudy (anonymous):

Actually it doesn't matter but Since the solution for this equation has to give 3 values of x.

OpenStudy (anonymous):

thanks!